[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Thu Jan 24 09:50:03 EST 2019

Logentry Text:
In addition to the restoration of the machine, MCC is also working on bringing back a few problematic cavities (the most problematic being 1L20) and replacing numerous circulator loads in the north in order to meet the set energy with an acceptable trip rate.  Unfortunately, this required work has resulted in a two day minimum delay to the start of the physics program.  Beam for physics is now expected for Friday, 2/1.  Therefore, shifts on Wednesday (1/30), Thursday (1/31), and Friday (2/1) owl are cancelled.  If all goes well, shifts will begin on day shift Friday (2/1).   

In addition, MCC reports that there exists two options moving forward to restore the machine to a reliable state.  One, they will need to spend a week helium processing on the problematic cavities (which is being recommended by SRF) which will ultimately delay delivery of physics beam by at least one week. Two, they will lower the energy and simply bypass the problematic cavities. The decision is being pushed up to management so that a decision can be made.

MCC will lock up the hall tonight at 5 pm during swing shift for DC power operations and beam line magnet testing.  The hall will re-open Thursday day shift. The earliest that we would lock up the hall for physics is now Wednesday (1/30) however, there are still numerous tasks to be completed in the hall prior to physics beam delivery.   If you require any assistance from Walter or the techs, please coordinate this closely with the RC prior to making the formal request so that we can determine the most efficient way to move forward.

Tasks to be completed in the Hall prior to lock up:

Aerogel installation (Vlad, ongoing)
HGC pump down, leak test, and installation (Brad, pump down complete, leak test ongoing, installation Friday 1/25)
Moller magnet cool down test (Dave G., ongoing)
Installation of lead shielding for J/Psi radiator (Techs, complete)
Beam line and scattering chamber pump down (Dave M., ongoing)
Target cool down with helium gas (Dave M., Friday 1/25) - contingent on beam line and scattering chamber pump down
Current load test of HMS & SHMS magnets (Steve L., Holly, Friday 1/25 ?)
Survey & alignment of downstream big beam line, harps, and target (S&A group, ?) 

Tasks to be completed at startup:

Jay has requested a beam deflection test whenever we start up in order to test the modifications to the Big BPMs.  He is requesting that we start at the smallest angle possible and follow the usual procedure.

Given the delay of physics beam, I feel it prudent that we cancel todays RC meeting and instead have our first RC meeting tomorrow (Friday, 1/25) at 11:00 pm in CH-200C (second floor of the counting house).  Note that we will typically meet at 1:00 pm however due to a scheduling conflict we need to meet a little early.

 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3645231
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190124/71f71f88/attachment.html>

More information about the Hallc_running mailing list